TGBT15A_915_146 - LBW bof file verification


* Verify the LBW BOF that will be used for pulsar modes. This will be done by actually passing it through the spectral line HPC backend.


  • Session begins at 11:00 AM
  • Will first test 100 MHz pulsar BOF (vpm_lbw_0100_2016_Apr_21_1156.bof) by running it in MODE7
  • Submitted CheckMode7 SB at 11:08
    • Scan is number 78
    • Scan ran without issue.
    • GBTIDL say all data are blanked, though the data tables seem to have non-zero values. Must be some issue with the switching signal.
    • Data are nans when opened with astropy.
  • Trying again but with switching set to tp_nocal, so there is only one phase and no blanking
    • Scan is number 79
    • No runtime issues
    • Data are nans again
  • Double checking original spectral line BOF file
    • Scan is number 80
    • No issues with this one
    • FITS state tables seem to be the same
  • Randy and Jason are checking the switching signal control in the new BOF file
    • The spec tick parameter is set to shortest possible resolution. This may be interfering with the daq software interpretation of the switching signals.
  • Took another scan (81) with spectral line BOF to check data display webpage.
  • Trying pulsar BOF but looking at the data display webpage (not sure where it intercepts the data...)
    • Scan is #82
    • Data still blanked in data display
  • Will test the 200 MHz pulsar BOF (vpm_lbw_0200_2016_Apr_21_1328.bof) by running it in MODE6
  • Submitted CheckMode6 SB
    • Scan is #83
    • No runtime issues, but data are still blanked
  • Break for lunch
  • Talked to Joe and Ray. Suggested checking INTEGRAT table. It is, indeed, all zeros. Will therefore set the "normalize_data" flag in vegas.conf to "false" to prevent dividing by INTEGRAT
  • Trying MODE6 again
    • "NORMALZD" keyword is zero in shmem, so it seems like the config file change was accepted
    • Scan is #84
    • No change
  • Will try MODE7 without normalization
    • Scan is #85
    • No change
  • Going to take two scans in each mode with the Archivist recording filtersnaps just in case we can play around with that data
    • Scans are number 86 and 87 (MODE7 and MODE6)
  • Returned vegas.conf to normal. Confirmed that BOF files and normalization flag are being read correctly with the production values.
  • Session ended at 12:53


  • Packet flow and balancing looked good with new BOF files, but for some reason the data is coming out nans. Will need to investigate switching signals and try again.
Topic revision: r1 - 2018-05-15, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback